Lorrimer Company has a job-order cost system. The following debits (credits) appeared in the Work-in-Process account for the month of June.

June 1, balance $12,000
June 30, direct materials 50,000

During the month of June, direct labor totaled $30,000 and $24,000 of overhead was applied to production. Finished Goods was debited $100,000 during June.

Lorrimer Company applies overhead at a predetermined rate of 80% of direct labor cost. Job number 83, the only job still in process at the end of June, has been charged with manufacturing overhead of $3,400. What was the amount of direct materials charged to Job number 83?

a.$3,400

b.$4,250

c.$7,580

d.$8,350
Answer #1

Calculate ending work in process :

Beginning work in process 12000
Direct material 50000
Direct labor 30000
Applied overhead 24000
Cost of goods manufactured -100000
Ending work in process 16000

Overhead applied = 3400

Direct labor = 3400*100/80 = 4250

Direct material = 16000-3400-4250 = 8350

So answer is d) $8350